Solve, using the method of separation of variables, the PDE

∂u

∂t +

∂u

∂x + 2e

tu = 0



1
Expert's answer
2022-01-12T18:40:35-0500

Since, given equation is unclear, we assume the given equation to solve is "\dfrac{∂u}{∂x} =2\dfrac{∂u}{∂t}+u=0 ; u(x,0)=6e^{-3x}"


Solution:

Given, "\dfrac{∂u}{∂x} =2\dfrac{∂u}{∂t}+u=0" ...(i)

Let the solution be "u(x,t)=X(x).T(t)" ...(ii)

Using (ii) in (i).

"X'T=2XT'+XT
\\ \Rightarrow (X'-X)T=2XT'
\\ \Rightarrow \dfrac{X'-X}X=\dfrac{2T'}T=k\ \ (say)"

Now, solving "\dfrac{X'-X}X=k"

"\Rightarrow \dfrac{X'}X=k+1"

On integrating,

"\log X=(k+1)x+\log C_1
\\ \Rightarrow \log X=(k+1)x\log e+\log C_1
\\ \Rightarrow \log X=\log e^{(k+1)x}+\log C_1
\\ \Rightarrow X=C_1e^{(k+1)x}"

On solving "\dfrac{2T'}T=k"

"\Rightarrow \dfrac{T'}T=\dfrac k2"

On integrating,

"\log T=\dfrac k2t+\log C_2
\\\Rightarrow T=C_2e^{kt/2}
\\ \therefore u(x,t)=X.T=C_1e^{(k+1)x}.C_2e^{kt/2}
\\=C_1C_2e^{(k+1)x+kt/2}"

When t=0,

"u(x,0)=C_1C_2e^{(k+1)x}=6e^{-3x}"

On comparing

"C_1C_2=6,k+1=-3\Rightarrow k=-4
\\ \therefore u(x,t)=6e^{-3x-2t}=6e^{-(3x+2t)}"